# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
To prove $(f_{1})$, we assume that $(V(P_{1})\cup V(P_{2}))-V(T)\neq\emptyset$. If some vertex $v$ in $V$ is contained in $V(T^{1})$, then $\alpha\in\overline{\varphi}_{1}(v)$ by $(b_{1})$ and $(d_{1})$. Since $P_{1}$ and $P_{2}$ both have some end in $V(T_{n+1}(x))$, by $(c_{1})$ and TAA, we have $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, a contradiction to our recent assumption. So both $P_{1}$ and $P_{2}$ have an end outside of $T^{1}$, and therefore the edge disjoint paths $P_{1}$ and $P_{2}$ give us (at least) two $\alpha$ colored boundary edges $f$ and $g$ of $T^{1}$ colored by $\alpha$ following $(c_{1})$, and $\alpha\notin\overline{\varphi}_{1}(T^{1})$. So by $(e_{1})$, one of $f$ and $g$ is incident to a vertex in $T^{1}$ other than $x$ and it is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Since $\alpha\notin\overline{\varphi}_{1}(T^{1})$ and $T^{1}$ is elementary by $(b_{1})$, we see that $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ such that $x_{2}$ is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Because $P_{z}(\pi,\alpha,\varphi_{1})$ and $T_{n+1}$ only share vertices on the subpath of $P_{z}(\pi,\alpha,\varphi_{1})$ from $z$ to $x$ (which is also $P_{z}(\pi,\alpha,\varphi)$) by the construction of $\varphi_{1}$ ($T_{n+1}$ is closed for $\pi$ and it only has one boundary edge colored by $\alpha$ which is the edge incident to $x$ under $\varphi_{1}$), we see that $x_{2}\notin V(T_{n+1})$ and the corresponding $(\pi,\alpha)$-exit path of $T^{1}$ under $\varphi_{1}$ does not share vertex with $T_{n+1}$. So $x\prec x_{2}$ along $T^{1}$ and $(f_{1})$ is established.
Now we have as claimed. If $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, we stop with $\mu=\varphi_{1}$ and $T=T^{1}$. We then have (I), (II), and $\mu$ is $(T_{j}(x)-x,D_{j},\varphi)$-stable by $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, $(b_{1})$ and $(a_{1})$. If $P_{1}$ and $P_{2}$ have different ends in $V$, then both ends of $P_{1}$ and $P_{2}$ in $V$ are contained in $T^{1}$, and therefore $T^{1}$ is not elementary by $(d_{1})$. Thus we have (III). Note that actually $P_{1}$ and $P_{2}$ can’t have different ends in this case, as otherwise we have a contradiction to $(b_{1})$. If $(V(P_{1})\cup V(P_{2}))-V(T^{1})\neq\emptyset$, we will in the following fashion recursively construct $\varphi_{i}$ with $i\geq 2$ from $\varphi_{i-1}$ and $x_{i}$ with an arbitrary color $\alpha_{i}\in\overline{\varphi}_{i-1}(x_{i})$, and have the following claim similarly as above with $T^{i}$ being a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$. Note that $\varphi_{1}$ also satisfies the claim below with $\alpha_{1}=\alpha$, $x_{1}=x$, $T^{0}=T_{j}$, and $\varphi_{0}=\varphi$.
@@
    $T^{i}$ is an elementary ETT with the same $j-1$-rungs under $\varphi_{i}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha_{i}$-colored edges of $P_{1}\cup P_{2}$ under $\varphi_{i}$ belong to $P_{z}(\pi,\alpha_{i},\varphi_{i})$ after the vertex $x_{i}$ along the direction from $z$.
@@
We will then prove the above claim inductively. By $(f_{k})$ for all $11\leq k\leq i-$ 1 and (1), we see that $x\prec x_{1},...,\prec x_{i}$ along $T^{i-1}$ and $x_{i}\notin T_{h}$. Following (1), $(a_{i-1})$, $(b_{i-1})$, $(c_{i-1})$, and $(d_{i-1})$, we have
(4) $\alpha_{i}\neq\alpha_{i-1}$ and $\pi,\alpha_{i}$, $\gamma_{h}$ (when $h\geq 1$), and $\alpha_{i-1}$ are all different, $\pi\notin S_{h}$ and $\alpha_{i}\notin\varphi_{i}(P_{1}\cup P_{2})$.
Clearly $\varphi^{\prime}$ is $(T^{i-1},D_{j-1},\varphi_{i-1})$-stable as $T^{i-1}$ is closed for both $\alpha_{i}$ and $\pi$ under $\varphi_{i-1}$ and $\varphi^{\prime}=\varphi_{i-1}/(G-T^{i-1},\alpha_{i},\pi)$. By $(f_{i-1})$, we see that $P_{x_{i}}(\alpha_{i},\alpha_{i-1},\varphi^{\prime})\cap T^{i-1}=\{x_{i}\}$. So $\varphi_{i}^{\prime}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable. As $T^{i-1}$ is closed for both $\alpha_{i}$ and $\pi$ under $\varphi_{i-1}$, all the $(\pi,\alpha_{i})$-chains intersect $T^{i-1}$ under $\varphi_{i}^{\prime}$ are contained in $G[V(T^{i-1})]$ except for $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ which contains $x_{i}$ who is the only $(\pi,\alpha_{i})$-exit of $T^{i-1}$ under $\varphi_{i}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$ are contained in $T^{i-1}$. Moreover, since $\alpha_{i}\neq\alpha_{i-1}$ by (4) and $x_{i}\prec x_{i+1}$ by $(f_{i-1})$, and $T^{i}$ is elementary by $(b_{i-1})$, we have $\alpha_{i}\notin(\overline{\varphi}_{i-1}(T^{i-1}(x_{i}))\cup\{\alpha_{i-1}\}$. So $\alpha_{i}\notin\varphi_{i-1}(P_{1}\cup P_{2})$ by $(c_{i-1})$. Thus $\pi$ is not used on edges of $P_{1}\cup P_{2}-G[V(T^{i-1})]$ and $\alpha_{i}$ is not used on edges of $(P_{1}\cup P_{2})\cap G[V(T^{i-1})]$ under $\varphi^{\prime}$ and $\varphi_{i}^{\prime}$. Therefore, since all vertices in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$ are contained in $T^{i-1}$, none of the $\alpha_{i}$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{i}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$. Consequently, we have $(e_{i})$ as we switch all the $\alpha_{i}$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ to $\pi$ by Kempe changes, and none of these $(\pi,\alpha_{i})$-chains intersects $T^{i-1}$ under $\varphi_{i}^{\prime}$. Furthermore,
@@
By $(a_{i})$ and (6), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{i}$ as under $\varphi_{i-1}$ if $h\geq 1$. By $(f_{k})$ for all $11\leq k\leq i-$ 1, we have $x_{i}\notin T_{j-1}$. Since $\varphi_{i}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable and $x_{i}\notin T_{j-1}$, we see that $\varphi_{i}$ is also $(T_{j-1},D_{j-1},\varphi_{i-1})$-stable ($T^{i-1}$ has the same rung under $\varphi_{i-1}$ as $T_{j}$ under $\varphi$ by $(b_{i-1})$), and thus is a $T_{j-1}+f_{j-1}$ mod $\varphi_{i-1}$ coloring by (A2.5). Therefore, $T^{i-1}(x_{i})$ is an ETT with the same $(j-1)$-rungs under $\varphi_{i}$ as $T^{i-1}$ under $\varphi_{i-1}$ following Corollary 3.1. Since $T^{i}$ is a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$ and $j-1\leq n$, we have $(b_{i})$ by (A1) on $T^{i}$ and $(b_{i-1})$.
The rest parts of the claim above are essentially the same as in the proof of the earlier claim with changed index, here we omit the details.
Now with the above claim, we see that we can always recursively construct $\varphi_{i}$ if $(V(P_{1})\cup V(P_{2}))-V(T^{i})\neq\emptyset$ with a new vertex $x_{i+1}$ with $x_{1}\prec x_{2},...,x_{i+1}$ along $T^{i}$ by $(f_{k})$ for all $1\leq k\leq i$ But since we have a finite graph, at a certain point we have to stop at some $\varphi_{i}$ with $(V(P_{1})\cup V(P_{2}))\subseteq T^{i}$. We denote the corresponding coloring $\varphi_{i}$ by $\mu$ and $T^{i}$ by $T$, respectively. Then (I) and (II) are satisfied by $\mu$ and $T$ following $(b_{i})$, and $\mu$ is $(T_{j}(x)-x,D_{j},\varphi)$-stable by $(a_{k})$ for all $1\leq k\leq i$. If $P_{1}$ and $P_{2}$ have different ends in $V$, then both ends of $P_{1}$ and $P_{2}$ in $V$ are contained in $T^{i}$, and therefore $T^{i}$ is not elementary by $(d_{i})$. Thus we have (III). Note that this is a contradiction to $(b_{i})$, and this contradiction is the key to the proof of Lemma 4.3.
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most $(n-1)$-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $RE$ finished with initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\alpha\in\overline{\varphi}(x)$ where $x$ is the $(\alpha,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$ and $x\notin V(T_{h})$. Suppose there’s an $(\alpha,\delta)$-ear $D$ of $T_{n+1}$ with one root $y$ such that $y\prec x$ along $T_{n+1}$. If $\tau\in\overline{\varphi}(T_{n+1}(x))$ with $\tau\notin S_{h}$ and $\eta$ is an arbitrary color where $\eta\notin\overline{\varphi}(T_{n+1}(x))$ and colors $\alpha,\eta,\delta$ and $\tau$ are all different, then there does not exist a path $P$ disjoint from $T_{n+1}$ with two ends $u$, $v$ and a vertex $w\in P$ with $w\notin\{u,v\}$ satisfying all of the following conditions:
@@
    All the $\eta$-colored edges of $P$ belong to a $(\eta,\tau)$-path $P_{1}=P_{r}(\eta,\tau,\varphi)$ containing the vertex $w\in P$ with one end $r\in T_{n+1}(x)$ and $\tau\in\overline{\varphi}(r)$, and all of these $\eta$-colored edges of $P$ are in the subpath $P_{wr^{\prime}}$ of $P_{1}$, where $r^{\prime}$ is the other end of $P_{1}$ and $P_{1}(wr^{\prime})$ does not intersect $T_{n+1}$.
@@
Let everything be as described as in Lemma 4.6, and assume otherwise there is a path $P$ disjoint from $T_{n}$ with two ends $u$, $w$ and a vertex $v\in P$ with $v\notin\{u,w\}$ satisfying conditions (i)-(iv). By the assumption in the statement of the current lemma, we have
@@
(4) There exits a color $\pi\in\overline{\varphi}(z)$ with $z\in V(T_{n+1}(x))$ such that $\pi,\alpha,\eta,\delta$,$\gamma_{h}$ (when $h\geq 1$) and $\tau$ are all different, and $\pi,\eta\notin S_{h}$.
@@
Similarly as in the proof of Lemma 4.3 (A3), we will construct a series of colorings $\varphi_{i}$. We first let $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)=\varphi/P_{ex}$, and then construct $\varphi_{1}$ from $\varphi_{1}^{\prime}$ by switching all the $\alpha$ colored edges of $P\cup P_{ex}\cup D$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes. Let $j$ be the smallest index such that $x\in T_{j}$. So $j-1\geq h$ by (1). Let $T^{1}$ be a closure of $T_{j}(x)$ under $\varphi_{1}$. We claim that $\varphi_{1}$ satisfies the following:
@@
    $T^{1}$ is an elementary ETT with the same $j-1$-rungs under $\varphi_{1}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha$-colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{1}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
To prove $(f_{1})$, we assume that $(V(P_{1})\cup V(P_{2}))-V(T^{1})\neq\emptyset$. If some vertex $v$ in $V$ is contained in $V(T^{1})$, then $\alpha\in\overline{\varphi}_{1}(v)$ by the fact that $V\cap V(T_{j}(x))=\emptyset$, $(a_{1})$, $(b_{1})$ and $(d_{1})$. Since both $P_{1}$ and $P_{2}$ have an end in $V(T_{n+1}(x))$, by $(c_{1})$ and TAA, we have $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, a contradiction to our recent assumption. So both $P_{1}$ and $P_{2}$ have an end outside of $T^{1}$, and therefore the edge disjoint paths $P_{1}$ and $P_{2}$ give us (at least) two $\alpha$ colored boundary edges $f$ and $g$ of $T^{1}$ colored by $\alpha$ following $(c_{1})$. Moreover, $\alpha\notin\overline{\varphi}_{1}(T^{1})$. So by $(e_{1})$, one of $f$ and $g$ is incident to a vertex in $T^{1}$ other than $x$ and it is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Since $\alpha\notin\overline{\varphi}_{1}(T^{1})$ and $T^{1}$ is elementary by $(b_{1})$, we see that $T^{1}$ has a $(\pi,\alpha)$-exit vertex $x_{2}$ in $P_{z}(\pi,\alpha,\varphi_{1})$ such that $x_{2}$ is after $x$ along $P_{z}(\pi,\alpha,\varphi_{1})$ from the direction starting at $z$. Because $P_{z}(\pi,\alpha,\varphi_{1})$ and $T_{n+1}$ only share vertices on the subpath of $P_{z}(\pi,\alpha,\varphi_{1})$ from $z$ to $x$ (which is also $P_{z}(\pi,\alpha,\varphi)$) by the construction of $\varphi_{1}$ ($T_{n+1}$ is closed for $\pi$ and it only has one boundary edge colored by $\alpha$ which is the edge incident to $x$ under $\varphi_{1}$), we see that $x_{2}\notin V(T_{n+1})$ and the corresponding $(\pi,\alpha)$-exit path of $T^{1}$ under $\varphi_{1}$ does not share vertex with $T_{n+1}$. So $x\prec x_{2}$ along $T^{1}$ and $(f_{1})$ is established.
Now we have as claimed. If $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, we stop with $\mu=\varphi_{1}$ and $T=T^{1}$. We then have (I), (II), and $\mu$ is $(T_{j}(x)-x,D_{j},\varphi)$-stable by $(V(P_{1})\cup V(P_{2}))\subseteq V(T^{1})$, $(b_{1})$ and $(a_{1})$. If $P_{1}$ and $P_{2}$ have different ends in $V$, then both ends of $P_{1}$ and $P_{2}$ in $V$ are contained in $T^{1}$, and therefore $T^{1}$ is not elementary by $(d_{1})$ and the fact that $V\cap V(T_{j}(x))=\emptyset$. Thus we have (III). Note that actually $P_{1}$ and $P_{2}$ can’t have different ends in this case, as otherwise we have a contradiction to $(b_{1})$. If $(V(P_{1})\cup V(P_{2}))-V(T^{1})\neq\emptyset$, we will in the following fashion recursively construct $\varphi_{i}$ with $i\geq 2$ from $\varphi_{i-1}$ and $x_{i}$ with an arbitrary color $\alpha_{i}\in\overline{\varphi}_{i-1}(x_{i})$, and have the following claim similarly as above with $T^{i}$ being a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$. Note that $\varphi_{1}$ also satisfies the claim below with $\alpha_{1}=\alpha$, $x_{1}=x$, $T^{0}=T_{j}$, and $\varphi_{0}=\varphi$.
@@
    $T^{i}$ is an elementary ETT with the same (j1)(j-1)-rungs under $\varphi_{i}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha_{i}$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{i}$ belong to $P_{z}(\pi,\alpha_{i},\varphi_{i})$ after the vertex $x_{i}$ along the direction from $z$.
@@
We will then prove the above claim inductively. By $(f_{k})$ for all $1\leq k\leq i-1$ and (1), we see that $x\prec x_{2}\prec...\prec x_{i}$ along $T^{i-1}$, and therefore $x_{i}\notin T_{h}\subseteq T_{j-1}$ by (1) and $(b_{i-1})$. Following (1), $(a_{i-1})$, $(b_{i-1})$, $(c_{i-1})$, and $(d_{i-1})$, we have
(4) $\alpha_{i}\neq\alpha_{i-1}$, $\pi,\alpha_{i}$, $\gamma_{h}$ (when $h\geq 1$), and $\alpha_{i-1}$ are all different, and $\pi\notin S_{h}$.
Clearly $\varphi^{\prime}$ is $(T^{i-1},D_{j-1},\varphi_{i-1})$-stable as $T^{i-1}$ is closed for both $\alpha_{i}$ and $\pi$ under $\varphi_{i-1}$ and $\varphi^{\prime}=\varphi_{i-1}/(G-T^{i-1},\alpha_{i},\pi)$. By $(f_{i-1})$, we see that $P_{x_{i}}(\alpha_{i},\alpha_{i-1},\varphi^{\prime})\cap T^{i-1}=\{x_{i}\}$. So $\varphi_{i}^{\prime}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable. As $T^{i-1}$ is closed for both $\alpha_{i}$ and $\pi$ under $\varphi_{i-1}$, all the $(\pi,\alpha_{i})$-chains intersecting $T^{i-1}$ under $\varphi_{i}^{\prime}$ are contained in $G[V(T^{i-1})]$ except for $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ which contains $x_{i}$ who is the only $(\pi,\alpha_{i})$-exit of $T^{i-1}$ under $\varphi_{i}^{\prime}$, and all vertices in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$ are contained in $T^{i-1}$. Moreover, since $\alpha_{i}\neq\alpha_{i-1}$ by (4) and $x_{i}\prec x_{i+1}$ by $(f_{i-1})$, and $T^{i}$ is elementary by $(b_{i-1})$, we have $\alpha_{i}\notin(\overline{\varphi}_{i-1}(T^{i-1}(x_{i}))\cup\{\alpha_{i-1}\}$. So $\alpha_{i}\notin\varphi_{i-1}(P_{1}\cup P_{2})$ by $(c_{i-1})$. Thus $\pi$ is not used on edges of $P_{1}\cup P_{2}-G[V(T^{i-1})]$ and $\alpha_{i}$ is not used on edges of $(P_{1}\cup P_{2})\cap G[V(T^{i-1})]$ under $\varphi^{\prime}$ and $\varphi_{i}^{\prime}$. Therefore, since all vertices in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$ are contained in $T^{i-1}$, none of the $\alpha_{i}$ colored edges of $P_{1}\cup P_{2}$ under $\varphi_{i}^{\prime}$ is in the subpath of $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ from $z$ to $x_{i}$. Consequently, we have $(e_{i})$ as we switch all the $\alpha_{i}$ colored edges of $P_{1}\cup P_{2}$ not in $P_{z}(\pi,\alpha_{i},\varphi_{i}^{\prime})$ to $\pi$ by Kempe changes, and none of these $(\pi,\alpha_{i})$-chains intersect $T^{i-1}$ under $\varphi_{i}^{\prime}$. Furthermore,
@@
By $(a_{i})$ and (6), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\varphi_{i}$ as under $\varphi_{i-1}$ if $h\geq 1$. By $(f_{k})$ for all $1\leq k\leq i-1$, we have $x_{i}\notin T_{j-1}$. Since $\varphi_{i}$ is $(T^{i-1}(x_{i})-x_{i},D_{j-1},\varphi_{i-1})$-stable and $x_{i}\notin T_{j-1}$, we see that $\varphi_{i}$ is also $(T_{j-1},D_{j-1},\varphi_{i-1})$-stable ($T^{i-1}$ has the same rung under $\varphi_{i-1}$ as $T_{j}$ under $\varphi$ by $(b_{i-1})$), and thus is a $T_{j-1}+f_{j-1}$ mod $\varphi_{i-1}$ coloring by (A2.5). Therefore, $T^{i-1}(x_{i})$ is an ETT with the same (j1)(j-1)-rungs under $\varphi_{i}$ as $T^{i-1}$ under $\varphi_{i-1}$ following Corollary 3.1. Since $T^{i}$ is a closure of $T^{i-1}(x_{i})$ under $\varphi_{i}$ and $j-1\leq n$, we have $(b_{i})$ by (A1) on $T^{i}$ and $(b_{i-1})$.
The rest parts of the claim above are essentially the same as in the proof of the earlier claim for $\varphi_{1}$ with changed index, here we omit the details.
Now with the above claim, we see that we can always recursively construct $\varphi_{i}$ if $(V(P_{1})\cup V(P_{2}))-V(T^{i})\neq\emptyset$ with a new vertex $x_{i+1}$ with $x_{1}\prec x_{2}\prec...\prec x_{i+1}$ along $T^{i}$ by $(f_{k})$ for all $1\leq k\leq i$ . But since we have a finite graph, at a certain point we have to stop at some $\varphi_{i}$ with $(V(P_{1})\cup V(P_{2}))\subseteq T^{i}$. We denote the corresponding coloring $\varphi_{i}$ by $\mu$ and $T^{i}$ by $T$, respectively. Then (I) and (II) are satisfied by $\mu$ and $T$ following $(b_{i})$, and $\mu$ is $(T_{j}(x)-x,D_{j},\varphi)$-stable by $(a_{k})$ for all $1\leq k\leq i$. If $P_{1}$ and $P_{2}$ have different ends in $V$, then both ends of $P_{1}$ and $P_{2}$ in $V$ are contained in $T^{i}$, and therefore $T^{i}$ is not elementary by $(d_{i})$ and the fact that $V\cap V(T_{j}(x))=\emptyset$. Thus we have (III). Note that this is a contradiction to $(b_{i})$, and this contradiction is the key to the proof of Lemma 4.3.
@@
Let $n$ be a nonnegative integer. Suppose (A1), (A2), (A2.5), (A3) hold for all ETTs with at most $n$ rungs, and (A3.5) holds for all ETTs with at most (n1)(n-1)-rungs. Let $T_{n+1}$ be a closed ETT with ladder $T_{0}\subset\dots\subset T_{n}\subset T_{n+1}$. Assume $T_{n+1}$ is $\operatorname{RE}$ finished with initial index $h$, $\delta$ is a defective color of $T_{n+1}$ under $\varphi$, and $\alpha\in\overline{\varphi}(x)$ where $x$ is the $(\alpha,\delta)$-exit vertex of $T_{n+1}$ with exit path $P_{ex}$ and $x\notin V(T_{h})$. Suppose there’s an $(\alpha,\delta)$-ear $D$ of $T_{n+1}$ with one root $y$ such that $y\prec x$ along $T_{n+1}$. If $\tau\in\overline{\varphi}(T_{n+1}(x))$ with $\tau\notin S_{h}$ and $\eta$ is an arbitrary color where $\eta\notin\overline{\varphi}(T_{n+1}(x))$ and colors $\alpha,\eta,\delta$ and $\tau$ are all different, then there does not exist a path $P$ disjoint from $T_{n+1}$ with two ends $u$, $v$ and a vertex $w\in V(P)$ with $w\notin\{u,v\}$ satisfying all of the following conditions:
@@
    All the $\eta$ colored edges of $P$ belong to an $(\eta,\tau)$-path $P_{1}=P_{r}(\eta,\tau,\varphi)$ containing the vertex $w\in V(P)$ with one end $r\in T_{n+1}(x)$ and $\tau\in\overline{\varphi}(r)$, and all of these $\eta$ colored edges of $P$ are in the subpath $P_{1}[w,r^{\prime}]$ of $P_{1}$, where $r^{\prime}$ is the other end of $P_{1}$ and $P_{1}[w,r^{\prime}]$ does not intersect $T_{n+1}$.
@@
Let everything be as described as in Lemma 4.6, and assume otherwise there is a path $P$ disjoint from $T_{n}$ with two ends $u$, $w$ and a vertex $v\in V(P)$ with $v\notin\{u,w\}$ satisfying conditions (i)-(iv). By the assumption in the statement of the current lemma, we have
@@
(4) There exits a color $\pi\in\overline{\varphi}(z)$ with $z\in V(T_{n+1}(x))$ such that $\pi,\alpha,\eta,\delta$$\gamma_{h}$ (when $h\geq 1$) and $\tau$ are all different, and $\pi,\eta\notin S_{h}$.
@@
Similarly as in the proof of Lemma 4.5, we will construct a series of colorings $\varphi_{i}$. We first let $\varphi_{1}^{\prime}=\varphi/P_{x}(\alpha,\delta,\varphi)=\varphi/P_{ex}$, and then construct $\varphi_{1}$ from $\varphi_{1}^{\prime}$ by switching all the $\alpha$ colored edges of $P\cup P_{ex}\cup D$ not in $P_{z}(\pi,\alpha,\varphi_{1}^{\prime})$ to $\pi$ by Kempe changes. Let $j$ be the smallest index such that $x\in T_{j}$. So $j-1\geq h$ by (1). Let $T^{1}$ be a closure of $T_{j}(x)$ under $\varphi_{1}$. We claim that $\varphi_{1}$ satisfies the following:
@@
    $T^{1}$ is an elementary ETT with the same j1j-1-rungs under $\varphi_{1}$ as $T_{j}$ under $\varphi$.
@@
    All the $\alpha$ colored edges of $P\cup P_{ex}\cup D$ under $\varphi_{1}$ belong to $P_{z}(\pi,\alpha,\varphi_{1})$ after the vertex $x$ along the direction from $z$.